# Rethinking the Learning Paradigm for Facial Expression Recognition

### v2

@@ IV-E Ablation Study
TABLE V: Comparing Softmax and vanilla PLL with different pre-training manners on FER without disambiguation module. “-PR” indicates pre-training.
@@
Maskfeat-PR + PLL 90.35 90.24 66.12
Figure 4: Test accuracy during fine-tuning on AffectNet-7.
IV-E1 Effectiveness of PLL Paradigm
To validate the PLL paradigm, we solely utilize the ResNet-18 model pre-trained on the FER dataset MS-Celeb-1M [67]. We directly compare the traditional Softmax loss with the vanilla PLL loss, employing only (eq. 1) without additional modules. The outcomes are reported in table V and fig. 4. (1) PLL loss consistently surpasses the Softmax loss, enhancing recognition accuracy by 2.13% on RAF-DB, 1.03% on FERPlus, and 0.63% on AffectNet-7. Thus, we experimentally verify that a good performance can be attained using ResNet-18 with the vanilla PLL loss. (2) As evident from fig. 4, during the initial training phases, pre-training establishes a solid foundation for feature initialization. However, as training advances, supervised learning gradually adjusts to ambiguous labels, making it sensitive to ambiguities within the label space, thus leading to performance degradation. In contrast, PLL exhibits greater resilience to the disruptive effects of ambiguous labels.
TABLE VI: The impact of pre-training (PR).
Method RAF FER Aff-7 Aff-8
Ours w/o PR 90.81 89.63 65.23 62.45
Ours 92.05 91.11 67.11 64.12
IV-E2 Effectiveness of Pre-training
As shown in table VI, we test the entire pipeline without using pre-training. Compared to “w/o PR”, pre-training has a significant performance gain. In this case, the robust features provided by pre-training further enhance the disambiguation ability of the PLL paradigm, even for vanilla PLL.
IV-E3 Feature-level Effects of Pre-trained SL, SSL
In contrast to supervised learning (SL), self-supervised learning (SSL) pre-training emerges as the optimal strategy for acquiring these resilient features in the FER task. In addition to employing ResNet-18 for SL pre-training, we also compare various other SSL pre-training methods (MoCo [69], SimMIM [48], MAE [49]) as presented in table VII. As previously highlighted, intuitively, the significance of unit-level information (e.g. eyes, nose, and mouth corners) surpasses that of pixel-level details (high-frequency nuances, lighting, contrast), because the structure of the face consists of multiple facial action units. This demonstrates that combining Maskfeat [50] with PLL forms a more effective self-supervised learning paradigm for FER. Notably, MoCo [69] demonstrates instead diminished effectiveness in FER due to its sensitivity to data augmentation, which has minimal influence on the facial structural components affected by variations in contrast and illumination within expression data.
TABLE VII: The impact of model architecture used for pre-training. No disambiguation module is used.
Method RAF FER Aff-7
ResNet-18 88.67 88.40 64.49
MoCo 88.59 88.14 63.98
MIM (SimMIM) 89.60 89.78 65.34
MIM (MAE) 89.99 89.70 65.79
MIM (Maskfeat) 90.35 90.24 66.12
Figure 5: The confidence correctness ratio of relabeled data on AffectNet-7.
IV-E4 Confidence-level performance
To intuitively observe and analyze the performance of PLL in terms of confidence, we artificially re-label and test 2k samples in the AffectNet-7 (the ratio of ambiguous labels is the highest in the FER dataset), and count the confidence ratio of these samples, as shown in fig. 5, compared to “Random init. + PLL”, the combination of features from “PLL” and “Pretrained” confers a greater capacity to resolve ambiguities. We argue that within this process, “PLL” plays a more pivotal role in terms of label disambiguation, yet its potential is fully harnessed when integrated with “Pretrained”.
IV-E5 Evaluation of Decoder and key modules
@@
To evaluate the effectiveness of the transformer decoder and other key modules, we conduct an ablation study on RAF-DB and FERPlus as shown in table VIII. From the results, based on the PLL paradigm, the transformer decoder improves accuracy by 0.94% on RAF-DB and 0.55% on AffectNet7. The results show that it uses the cross-attention module to obtain the correlation of features and category labels via learnable label embedding, which in turn provides a more adequate foundation for the PLL disambiguation.
TABLE VIII: The impact of decoder, $\mathcal{L}_{uniform}$, and revision confidence in our disambiguation module.
Decoder $\mathcal{L}_{uniform}$ Re-conf RAF Aff-7
@@
IV-E6 Ablation for Threshold of Revision Confidence
The revision confidence module relies on the output of the transformer decoder. In addition, the disambiguation ability of PLL is determined by setting a threshold for the uncertainty of the candidate set. We aim to find and revise the sample, whose confidence is less than the threshold. To evaluate the influence of different thresholds for disambiguation. We conduct an ablation study as fig. 6 and fig. 7 show.
Figure 6: The accuracy (%) with different thresholds on RAF-DB and FERPlus.
Figure 7: The accuracy (%) with a different thresholds on AffectNet7 and AffectNet8.
IV-E7 Influence of parameter K
We conducted the influence of K in different settings. It is somewhat sensitive, and we have set k=2 for optimal performance. After conducting additional experiments, we observed that in 95% of cases, a true label is included for k=2, and this increases to 99% for k=3. In terms of performance, on the RAF-DB dataset, the recognition rates are 92.35% for k=2, 91.94% for k=3, and 90.01% for k=4. It is evident that k is a bit sensitive to different settings, which is possibly due to the trade-off between increasing the possibility of including potential ground truth labels and the simultaneous increase in uncertainty and noise. Consequently, as k increases, the performance gradually declines.
IV-E8 Ablation for Temperature Hyperparameter $\tau$ and Calculation of the Confusion Matrix
To correct the effects of unbalanced data classes, we propose $\mathcal{L}_{uniform}$, whose aim is to distribute the query embedding uniformly over the hypersphere to help the decoder better find the correlation between features and category labels. Inspired by [70], a small temperature coefficient is beneficial to distinguish the embedding of one category from the embedding of other categories. To verify its validity, we conducted an ablation study, as table IX shows.
TABLE IX: The accuracy (%) with different $\tau>0$ on RAF-DB and FERPlus.
@@ IV-E8 Ablation for Temperature Hyperparameter $\tau$ and Calculation of the Confusion Matrix
Figure 8: The confusion matrix of the RAF-DB. Left part without $\mathcal{L}_{uniform}$, right part with $\mathcal{L}_{uniform}$.
Figure 9: The confusion matrix of the FERPlus. Left part without $\mathcal{L}_{uniform}$, right part with $\mathcal{L}_{uniform}$.
To demonstrate the validity of $\mathcal{L}_{uniform}$ more intuitively, we calculated the confusion matrix for each category and calculated overall We show the results for the RAF-DB and FERPlus confusion matrix with respect to $\mathcal{L}_{uniform}$ in fig. 8, fig. 9, respectively.
@@
From the confusion matrix on RAF-DB and FERPlus, we improve average accuracy on RAF-DB and FERPlus by 1.4%, 5.37%. In the head class, such as the ”HAPPY” in RAF-DB, more test samples are classified to the tail class after using the $\mathcal{L}_{uniform}$ constraint, which is beneficial to side effects of class imbalance and helps to eliminate bias to head class in the PLL disambiguation process; for FERPlus, the average accuracy improvement is relatively large, since its test samples are in the tail class, such as contempt, with only 15 images, which is an order of magnitude difference compared to the head class. When calculating the average accuracy, the improvement in accuracy in the tail class contributes significantly to the overall average precision.
IV-F Visualization of PLL disambiguation
To illustrate the disambiguation process of PLL more visually, we show the more visualization results of PLL disambiguation on RAF-DB, FERPlus, and AffectNet7 in fig. 10, fig. 11, and fig. 12, respectively.
Figure 10: The grey part of the figure shows the GT labels corresponding to the samples in FERPlus; the blue part shows the candidate set we constructed for the samples, where the strikethroughs mark the results of PLL disambiguation for error removal. The predicted labels are consistent with our intuition.
Figure 11: The grey part of the figure shows the converted label, equivalent to GT, which is selected by the annotators from the crowdsourcing results in FERPlus. The blue part shows the candidate set we constructed for the samples corresponding to the converted(GT) labels, where the strikethroughs mark the results of PLL disambiguation for error removal. The predicted labels are consistent with our intuition.
Figure 12: The grey part of the figure shows the converted label, equivalent to GT, which is selected by the annotators from the crowdsourcing results in FERPlus. The blue part shows the candidate set we constructed for the samples corresponding to the converted(GT) labels, where the strikethroughs mark the results of PLL disambiguation for error removal. The predicted labels are consistent with our intuition.
IV-G Time/Memory Consumption of Proposed Method
We trained 600 epochs using 4 32GB Tesla V100s, requiring 3 days for pre-training, which takes longer and relatively more memory resources. The related pre-trained checkpoint will be released to help with FER downstream tasks. As for fine-tuning, we do an ablation on RAF-DB with batch size=72/128/256 and it takes around 1.43/1.2/1.18 hours and around 11/16/30 GB of memory, which is similar to DMUE[29])/MA-NET[56]/EAC[59]. However, we use similar resources to achieve better performance.
V Conclusion

### v4

@@ 4.4. Ablation Study
Table 5. Comparing Softmax and vanilla PLL with different pre-training manners on FER without disambiguation module. “-PR” indicates pre-training.
@@
[…]
4.4.1. Effectiveness of PLL Paradigm
To verify the effectiveness of our PLL paradigm, we only use the ResNet-18 pre-trained on FER dataset MS-Celeb-1M (guo2016ms) and pre-trained Maskfeat as the backbone. we compare the traditional Softmax Loss to Vanilla PLL Loss only using Eq.(2) without any other modules and report the results in Table 5. (1) Compared with Softmax loss, PLL loss exceeds it across the board, improving the recognition accuracy by 2.13% on RAF-DB, 1.03% on FERPlus, and 0.63% on AffectNet-7 in terms of ResNet-18. We experimentally verified that a good performance can be achieved even if only using ResNet-18 with the vanilla PLL loss. The “Maskfeat-PR” setting yields a similar conclusion. (2) Table 4 reveals that most ResNet-based methods employ supervised learning to represent the label space distribution for samples, which is susceptible to the influence of incorrect labels in the label space. From the results, simplifying FER as a supervised learning task is not optimal.
4.4.2. Feature-level Effects of Pre-trained SL, SSL
Table 6. Different pre-training manners with vanilla PLL on FER without disambiguation module. “-PR” indicates pre-training.
Method RAF-DB FERPlus AffectNet-7
ResNet-18 + PLL 88.67 88.40 64.49
MoCo + PLL 88.59 88.14 63.98
SimMIM-PR + PLL 89.60 89.78 65.34
MAE-PR + PLL 89.99 89.70 65.79
[…]
Figure 4. Test accuracy during fine-tuning on AffectNet-7.
While vanilla PLL is effective, we demonstrate that robust features can significantly enhance the ability to disambiguate the PLL paradigm. We argue that self-supervised pre-training is the optimal method for obtaining robust features in the FER task. Thus, we also compare other SSL pre-training approaches (MoCo (he2020momentum), SimMIM (xie2022simmim), MAE (he2022masked)) as shown in Table 6, and demonstrate that combining Maskfeat (wei2022masked) with PLL is a more effective self-supervised learning paradigm for FER. We note that the MoCo (he2020momentum), which is sensitive to data augmentation, is not very effective in the FER because the variations in contrast and illumination in the expression data have little effect on the facial structural units. As shown in Figure 4, we conduct a comparison between “Pretrained MIM + PLL”, “Pretrained MIM + SoftMax” (which denotes a pre-trained encoder with a different loss function), and “Random Initialization + SoftMax” (which denotes random initialization of the encoder. (1) “Random Initialization + SoftMax” easily suffers from incorrect labels in the dataset during the whole fine-tuning, which further impedes the model from learning the high-quality features, thereby undermining the model’s classification ability and decreasing its robustness. Compared to it, “Pretrained MIM + SoftMax” slows down overfitting to incorrect labels at the early stage, which is provably robust against the incorrect labels during fine-tuning. (2) At the initial stage of fine-tuning, “Pretrained MIM + SoftMax” also achieves superior performance. However, as training goes on, the interference of incorrect labels gradually increases, and the model begins to overfit these incorrect labels, ultimately resulting in performance degradation. Whereas, at the later stage of fine-tuning, “Pretrained MIM + PLL” resists the interference caused by incorrect labels, leading to robust performance in FER.
Figure 5. The confidence correctness ratio of relabeled data on AffectNet-7.
To quantitatively and intuitively observe and analyze the performance of PLL in terms of confidence, we artificially re-labeled and tested 2k samples in the AffectNet-7 (the ratio of incorrect labels is the highest in the FER dataset), and counted the confidence ratio of these samples, as shown in the Figure 5, during the fine-tuning, compared to “Random initialization + PLL”, “Pretrained MIM + PLL” disambiguates these samples more effectively by combining the robust features obtained by SSL pre-training. They complement each other, but PLL plays a more important role in the disambiguation process. We have attached some PLL disambiguation results as shown in  Figure 7 on FERPlus. The experiment results demonstrate the effectiveness of the PLL paradigm. For RAF-DB, and AffectNet7, see Figure 8, and Figure 9, respectively.
4.4.3. Evaluation of Decoder and key modules
@@
To evaluate the effectiveness of the transformer decoder and other key modules, we conduct an ablation study on RAF-DB and FERPlus as shown in Table 7. From the results, based on the PLL paradigm, the transformer decoder improve accuracy by 0.94% on RAF-DB and 0.55% on AffectNet7. The results show that it uses the cross-attention module to obtain the correlation of features and category labels via learnable label embedding, which in turn provides a more adequate foundation for the PLL disambiguation.
Table 7. Evaluation of transformer decoder-based label disambiguation module with PLL. Accuracy (%) and gains over the baseline are reported. “Decoder” denotes the transformer decoder, label embedding regularization is $\mathcal{L}_{uniform}$”, and revision confidence is “Re-conf”.
Decoder $\mathcal{L}_{uniform}$ Re-conf RAF-DB Gain AffectNet-7 Gain
@@
4.4.4. Ablation for Threshold of Revision Confidence
The revision confidence module relies on the output of the transformer decoder. In addition, the disambiguation ability of PLL is determined by setting a threshold for the uncertainty of the candidate set. We aim to find and revise the sample, whose confidence is less than the threshold. To evaluate the influence of different thresholds for disambiguation. We conduct an ablation study as fig. 6(a) and fig. 6(b) show.
(a) The accuracy (%) with different thresholds on RAF-DB and FERPlus.
(b) The accuracy (%) with different thresholds on AffectNet7 and AffectNet8.
Figure 6. Accuracy with different thresholds on multiple datasets.
4.4.5. Ablation for Temperature Hyperparameter $\tau$ and Calculation of the Confusion Matrix
To correct the effects of unbalanced data classes, we propose $\mathcal{L}_{uniform}$, whose aim is to distribute the query embedding uniformly over the hypersphere to help the decoder better find the correlation between features and category labels. Inspired by (wang2021understanding), a small temperature coefficient is beneficial to distinguish the embedding of one category from the embedding of other categories. To verify its validity, we conducted an ablation study, as table 8 shows.
Table 8. The accuracy (%) with different $\tau>0$ on RAF-DB and FERPlus.
@@ 4.4.5. Ablation for Temperature Hyperparameter $\tau$ and Calculation of the Confusion Matrix
Figure 7. The grey part of the figure shows the converted label, equivalent to GT, which is selected by the annotators from the crowdsourcing results in FERPlus. The blue part shows the candidate set we constructed for the samples corresponding to the converted (GT) labels, where the strikethroughs mark the results of PLL disambiguation for error removal. The predicted labels are consistent with our intuition.
Figure 8. The grey part of the figure shows the converted label, equivalent to GT, which is selected by the annotators from the crowdsourcing results in FERPlus. The blue part shows the candidate set we constructed for the samples corresponding to the converted(GT) labels, where the strikethroughs mark the results of PLL disambiguation for error removal. The predicted labels are consistent with our intuition.
Figure 9. The grey part of the figure shows the converted label, equivalent to GT, which is selected by the annotators from the crowdsourcing results in FERPlus. The blue part shows the candidate set we constructed for the samples corresponding to the converted(GT) labels, where the strikethroughs mark the results of PLL disambiguation for error removal. The predicted labels are consistent with our intuition.
4.5. Influence of parameter K
We conducted the influence of K in different settings. It is somewhat sensitive, and we have set k=2 for optimal performance. After conducting additional experiments, we observed that in 95% of cases, a true label is included for k=2, and this increases to 99% for k=3. In terms of performance, on the RAF-DB dataset, the recognition rates are 92.05% for k=2, 91.94% for k=3, and 90.01% for k=4. It is evident that k is a bit sensitive to different settings, which is possibly due to the trade-off between increasing the possibility of including potential ground truth labels and the simultaneous increase in uncertainty and noise. Consequently, as k increases, the performance gradually declines.
Figure 10. The confusion matrix of the RAF-DB. Left part without $\mathcal{L}_{uniform}$, right part with $\mathcal{L}_{uniform}$.
Figure 11. The confusion matrix of the FERPlus. Left part without $\mathcal{L}_{uniform}$, right part with $\mathcal{L}_{uniform}$.
To demonstrate the validity of $\mathcal{L}_{uniform}$ more intuitively, we calculated the confusion matrix for each category and calculated overall We show the results for the RAF-DB and FERPlus confusion matrix with respect to $\mathcal{L}_{uniform}$ in Figure 10, Figure 11, respectively.
@@ 4.5. Influence of parameter K
From the confusion matrix on RAF-DB and FERPlus, we improve average accuracy on RAF-DB and FERPlus by 1.4%, 5.37%. In the head class, such as the ”HAPPY” in RAF-DB, more test samples are classified to the tail class after using the $\mathcal{L}_{uniform}$ constraint, which is beneficial to side effects of class imbalance and helps to eliminate bias to head class in the PLL disambiguation process; for FERPlus, the average accuracy improvement is relatively large, due to the fact that its test samples are in the tail class, such as contempt, with only 15 images, which is an order of magnitude difference compared to the head class. When calculating the average accuracy, the improvement in accuracy in the tail class contributes significantly to the overall average precision.
5. Conclusion
@@
6. Safe and Responsible Innovation Statement
This work aims to improve the accuracy and robustness of facial expression recognition systems. We use ethically sourced and consented datasets, ensuring no sensitive or personally identifiable information is involved. To mitigate potential biases, we consider demographic diversity during model training and evaluation. As facial expression recognition can be misused for surveillance or behavioral manipulation, we emphasize that its deployment must comply with ethical guidelines and legal standards, promoting human-centered and responsible applications.